OpenStudy (anonymous):

tan70^o=2tan50^o+tan20^o prove this?

hartnn (hartnn):

used this formula : \(tan (A-B)=\frac{tanA-tanB}{1+tanAtanB}\)

hartnn (hartnn):

tan(70-20)=tan50 \(\frac{(tan70-tan20)}{(1+tan20 tan70)}\) =tan50 tan20 tan70=cot70 tan70=1 (tan70-tan20)/2=tan50 tan70-tan20=2tan50 tan70=tan20+2tan50